### v3

@@ VI.1 Temperature
$4\pi T$ $=$ $\frac{d}{dr}f(r)^{2}=2\frac{r_{+}}{l^{2}}+\frac{(d-2n-1)}{nr_{+}}\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)$ (55)
-- $\frac{1}{n!}\frac{\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}}{\left(1-e^{-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)}\left(\frac{r_{+}}{a}\right)^{\frac{n-1}{n}(d-2n-1)}$
To analyze the expression above it is worth to recall the presence of the two length scales $a$ and $l$. One can notice that the two first terms of Eq.(55) not only are independent of $a$, but coincide with the expression of the temperature found in Aros et al. (2001); Crisostomo et al. (2000) for the vacuum LnFDGS AdS solution. Therefore, to address the difference, one must play attention to the last term. In this connection, one can notice that this term dependents only on $r_{+}/a$. The structure of the third term for $r_{+}\gg a$ has the form
$\lim_{r_{+}/a\rightarrow\infty}\frac{\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}}{\left(1-e^{-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)}\left(\frac{r_{+}}{a}\right)^{\frac{n-1}{n}(d-2n-1)}\approx 0,$ (56)
Thus, we analytically observe that in a scenario where $a\ll r_{+}$, the temperature of the vacuum LnFDGS AdS solution becomes indistinguishable from our RBH LnFDGS AdS solution.
@@
[…]
@@
From Figure 2, we can check that the sign of the derivative $\frac{dM}{dr_{+}}$ is always positive, i.e., $\frac{dM}{dr_{+}}\geq 0$. Thus, the sign of the derivative $\frac{dT}{dM}=\left(\frac{dT}{dr_{+}}\right)\left(\frac{dM}{dr_{+}}\right)^{-1}$ depends only on the sign of the derivative $\frac{dT}{dr_{+}}$. In summary, for our solution, the temperature exhibits three regions from left to right: i) $\frac{dT}{dM}>0$ due to the action of matter sources at small scales, ii) and iii) which coincide with the vacuum LnFDGS AdS case, where $\frac{dT}{dM}<0$ and $\frac{dT}{dM}>0$, respectively.
Figure 13: The temperature for $d=6$, $n=2$, $a=1$ and $l=10$
VI.2 Heat Capacity
Given that there is just a single parameter in this case, says $r_{+}$, is the heat capacity can be written as
$C=\frac{dE}{dT}=\left(\frac{dT}{dM}\right)^{-1}=\frac{dM}{dr_{+}}\left(\frac{dT}{dr_{+}}\right)^{-1}$ (57)
As mentioned before, from Figure 2, the derivative $\frac{dM}{dr_{+}}$ is always positive, i.e., $\frac{dM}{dr_{+}}\geq 0$. Since the sign of the derivative $\frac{dT}{dM}=\left(\frac{dT}{dr_{+}}\right)\left(\frac{dM}{dr_{+}}\right)^{-1}$ depends only on the sign of the derivative $\frac{dT}{dr_{+}}$, consequently, the sign of the heat capacity also depends solely on this last derivative.
Before to compute the heat capacity, it is worth to recall the heat capacity of the vacuum LnFDGS AdS case Crisostomo et al. (2000); Aros et al. (2001). This is given by
$C_{\text{vac}}=\frac{\left(\frac{\gamma(d-2n-1)}{2r_{+}}+(d-1)\frac{r_{+}}{2l^{2}}\right)\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)^{n-1}r_{+}^{d-2n-1}}{\left(\frac{2}{l^{2}}-\frac{(d-2n-1)}{nr_{+}^{2}}\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)+\frac{2(d-2n-1)}{nl^{2}}\right)}$ (58)
One can notice $C_{\text{vac}}$ has a divergency. This defines the phase transition noticed in Crisostomo et al. (2000).
To compute the heat capacity in this case one only need to consider the mass function Eq.(54) and the temperature (55) above.
$C$ $=$ $\frac{\left(\frac{\gamma(d-2n-1)}{2r_{+}}+(d-1)\frac{r_{+}}{2l^{2}}\right)\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)^{n-1}r_{+}^{d-2n-1}}{\left(\frac{2}{l^{2}}-\frac{(d-2n-1)}{nr_{+}^{2}}\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)+\frac{2(d-2n-1)}{nl^{2}}\right)}\left(1-\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)^{-1}$ (59)
-- $\frac{n\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)^{n}r_{+}^{d-2n-1}}{\left(\frac{2}{l^{2}}-\frac{(d-2n-1)}{nr^{2}}\left(\gamma+\frac{r^{2}}{l^{2}}\right)+\frac{2(d-2n-1)}{nl^{2}}\right)}\times n\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}\left(\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)^{n-1}$
$\left(1-\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)^{-2}\left(\frac{d-1}{n}\right)\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}-1}$
This, in turn, allows one to analytically conclude that
$\lim_{r_{+}/a\rightarrow\infty}C=C_{\text{vac}}$ (60)
This is quite interesting in two aspects. Firstly, due to $r_{+}/a\gg 1$, as mentioned earlier, it is a natural consideration for solar mass black holes, and therefore the standard thermodynamics must be recovered in that regime. Secondly, it must be recalled that, unlike for non-regular black holes, the radius of the outer horizon is bounded by $r_{ext}<r_{+}$ and thus cannot be arbitrarily small. In this way, the limit of interest is $r_{ext}\rightarrow r_{+}$, which represents the limit as $T\rightarrow 0$.
VI.2.1 On the radial evolution
A usual interpretation of the sign of the heat capacity is as follows Estrada and Aros (2023, 2025): a positive heat capacity indicates that when the temperature decreases (increases), the black hole emits (absorbs) thermal energy, thus $dM<0$ ($dM>0$) in the black hole, in order to reach thermodynamic equilibrium with the external environment, i.e., the black hole is stable. Conversely, a negative heat capacity implies that, if the temperature increases (decreases), the black hole also emits (absorbs) thermal energy toward the external environment, i.e., the black hole is unstable. As we will see below, under this interpretation, we can assume that at Planck scales, our matter sources induce a phase transition such that the black hole becomes stable at these scales, leading to a black hole remnant rather than a divergent temperature.
@@ VI.2.1 On the radial evolution
On the other hand, related to the analysis of phase transitions, this can be easily unfolded by using the fact that $C=\left(\frac{dT}{dM}\right)^{-1}\sim\left(\frac{dT}{dr_{+}}\right)^{-1}$ in graph 13. Moving from right to left, we can notice that once the vacuum LnFDGS AdS solution and our RBH LnFDGS AdS solution are no longer indistinguishable, the heat capacities of both solutions also differ. While the heat capacity of the vacuum AdS solution remains negative, meaning that its complete evaporation, as the parameter $M$ decreases to $M\to 0$, would imply the temperature having to rise to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until the black hole remnant is reached at $T=0$, where the evaporation process halts. As mentioned earlier, this could occur at scales close to the Planck scale. That is, the nature of our matter sources, along with avoiding the formation of a singularity, would lead to the formation of a remnant. Thus, in our solution, there would be two phase transitions, from right to left: i) $C>0$ where $dT/dM>0$, ii) $C<0$ where $dT/dM<0$ (these two coincide with the vacuum LnFDGS AdS case, where both solutions are indistinguishable), iii) $C>0$ where $dT/dM>0$.
@@ VII DISCUSSION AND SUMMARIZE
Motivated by the following facts:
    The recent higher-dimensional model to suppress physical singularities from Estrada and Aros (2025) for Pure Lovelock gravity, where the energy density encodes the gravitational information of the vacuum solution through the Kretschmann scalar.
    The theoretical importance of incorporating a negative cosmological constant in various physical problems such as the AdS/CFT correspondence, the thermodynamic interpretation of the negative cosmological constant as a positive pressure in the universe, etc.
    The invalidity of including a negative cosmological constant in PL theories in the spherically symmetric case, since its inclusion leads to the appearance of a curvature singularity. This singularity is such that, although the metric tensor is regular for a radius greater than the event horizon, the Ricci and Kretschmann invariants diverge at a point within this location.
    It is worth mentioning that directly relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution in the same way as for PL theory is complicated, because several additional terms would appear in the gravitational tension.
    Related to the study of shadows, recently in reference Paithankar and Kolekar (2023), relationships were found between the event horizon, the photon sphere radius, and the shadow size for the vacuum solution without a cosmological constant in Pure Lovelock. However, the inclusion of matter and the cosmological constant makes it impossible to find analytical relationships for our case study of LnFDGS AdS. Therefore, it is useful to propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior.
To address the points described above, instead of considering the PL action, which takes into account only one term in the Lovelock series, we use the definition of the coupling constants from references Crisostomo et al. (2000); Aros et al. (2001), such that, by considering all the terms in the Lovelock series up to order $n=[d/2]$, the equations of motion possess an $n$-fold degenerate ground state AdS. In this way, the inclusion of the negative cosmological constant does not lead to the appearance of a curvature singularity.
@@
The obtained solution is such that there exists a value of the radial coordinate $r_{-}<r=r_{*}<r_{+}$, greater than the inner horizon and less than the event horizon, such that for radial values slightly greater than $r_{*}$, the vacuum and singular LnFDGS solution from references Crisostomo et al. (2000); Aros et al. (2001) becomes indistinguishable from our regular LnFDGS solution. In the case that $r_{*}$ is of the order of the Planck scales, the geometric differences between both solutions, and specifically, the suppression of the singularity, would occur at quantum scales.
Related to the study of shadows, a method has been proposed to obtain numerical graphical relationships between the event horizon, the photon sphere radius, and the shadow size in our case study. The case $n=2,d=7$ has been used as an example to explain our procedure. However, it is straightforward to check that our proposal is applicable to other values of $n$ and $d$. We have also displayed the graphical behavior for the cases $n=2,d=6,7,8$ and $n=3,d=8,9,10$. We can note that, if we use a common event horizon value as a reference for the same value of $n$, the photon sphere radius decreases as the number of dimensions increases. Additionally, if we use a common photon sphere radius value as a reference for the same value of $n$, the size of the black hole shadow radius tends to decrease as the number of dimensions increases. It is also important to mention that in all the cases studied numerically, the condition $r_{+}<r_{sp}<r_{sh}$ is satisfied.
We have also provided a speculative way to compare theoretical results obtained for Lovelock gravity with the experimental results from the EHT for M87. To this end, we have used the vacuum gravitational potential of the Lovelock solution as a comparison parameter, such that our Equation (53) reduces to Constraint (50) for $n=1,\ d=4$, which corresponds to the data provided by the EHT for M87. We tested whether the vacuum gravitational potential yields values consistent with the constraints given by Equation (53). Graphically, we have found that, depending on the chosen parameters, the theoretical results obtained for Lovelock gravity can match the experimental data. In particular, we find that our results agree with the experimental results from the EHT for M87 in our plots corresponding to $n=2,\ d=7,8$ and $n=3,\ d=10,11$. It seems that, for a fixed value of $n$, the parameters tend to better match the experimental results as the number of dimensions increases.
@@
Regarding the analysis of heat capacity and radial evolution, moving from right to left in graph 13, we can notice that once the vacuum LnFDGS AdS solution and our RBH LnFDGS AdS solution are no longer indistinguishable, the heat capacities of both solutions also differ. While the heat capacity of the vacuum AdS solution remains negative, meaning that its complete evaporation, as the parameter $M$ decreases to $M\to 0$, would imply the temperature rising to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until the black hole remnant is reached at $T=0$, where the evaporation process halts. We could assume that at Planck scales, our matter sources induce a phase transition such that the black hole becomes stable at these scales, leading to a black hole remnant rather than a divergent temperature.
appendices
Appendix A Hypergeometric Confluent Function Expression
The confluent hypergeometric function above can be written as
${}_{1}F_{1}\left(n,n+1,-y\right)=\frac{n!}{y^{n}}\left(1-e^{-y}\sum_{i=0}^{n-1}\frac{y^{i}}{i!}\right)$ (61)
and thus in the case at hand
${}_{1}F_{1}\left(n,n+1,-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)=n!\left(\frac{a}{r}\right)^{d-1}\left(1-\exp{\left(-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r}{a}\right)^{\frac{i}{n}(d-1)}\right)$ (62)
The derivative of this function is given by
$\frac{d}{dy}\left({}_{1}F_{1}\left(n,n+1,-y\right)\right)=\frac{n}{y}\left(e^{-y}-{}_{1}F_{1}\left(n,n+1,-y\right)\right)$ (63)
@@ Acknowledgements
This work of RA was partially funded through FONDECYT-Chile 1220335. Milko Estrada is funded by the FONDECYT IniciaciĂłn Grant 11230247.

### v4

@@ VI.1 Temperature
$4\pi T$ $=$ $\frac{d}{dr}f(r)\Big|_{r=r_{+}}=2\frac{r_{+}}{l^{2}}+\frac{(d-2n-1)}{nr_{+}}\left(\gamma+\frac{r_{+}^{2}}{l^{2}}\right)$ (51)
- $\frac{1}{n!}\frac{\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}}{\left(1-e^{-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}}\displaystyle\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)}\left(\frac{r_{+}}{a}\right)^{\frac{n-1}{n}(d-2n-1)}$
In Eq. (51), we can notice that the first two terms are independent of $a$, and furthermore, they match the expression for the temperature found in Aros et al. (2001); Crisostomo et al. (2000) for the vacuum LnFDGS AdS solution. Regarding the third term in Eq. (51), we can analytically observe that for a scenario where $r_{+}\gg a$, condition
$\lim_{r_{+}/a\rightarrow\infty}\frac{\exp{\left(-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}\right)}}{\left(1-e^{-\left(\frac{r_{+}}{a}\right)^{\frac{d-1}{n}}}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r_{+}}{a}\right)^{\frac{i}{n}(d-1)}\right)}\left(\frac{r_{+}}{a}\right)^{\frac{n-1}{n}(d-2n-1)}\approx 0$ (52)
holds. Thus, we find analytically that in a regime where $a\ll r_{+}$, the temperature of the vacuum LnFDGS AdS solution becomes indistinguishable from that of our RBH LnFDGS AdS solution. This regime could be of physical interest in the case where $a$ is of the order of the Planck scale, while $r_{+}$ corresponds to a realistic radius—on the order of the radius of the rotating black hole observed in LIGO GW150914, $r_{+}\sim 103,\text{km}$—which is much larger than the Planck scale.
@@
In Fig. 11, the generic numerical behavior of the temperature as a function of the event horizon $r_{+}$ is displayed. We can observe that this behavior is generic for other values of the parameters involved. In this figure, the horizontal axis can be considered to be of the order of Planck units, $\ell_{p}$. Two curves are shown. The blue curve represents the temperature of the vacuum LnFDGS AdS solution Aros et al. (2001); Crisostomo et al. (2000) as a function of the event horizon. The orange curve corresponds to the temperature of our RBH LnFDGS AdS solution. It can be observed that there exists a finite value slightly greater than the extremal radius, $r_{*}>r_{ext}$, where the temperatures of both cases become indistinguishable.
It is of interest to discuss what happens for values smaller than $r_{*}$, that is, radial values prior (from left to right) to the point where both temperatures become indistinguishable:
  • The temperature of the vacuum solution increases and hypothetically diverges in the limit $r_{+}\to 0$.
  • On the other hand, in the case of the regular solution, we can note that the temperature decreases until it vanishes for the value $M=M_{\text{min}}$, where $r_{-}=r_{+}=r_{\text{ext}}$.
In this sense, it is well known that the point of zero temperature is associated with a black hole remnant, what remains once evaporation ceases. In this context, in Figure 11 below, we observe that a remnant indeed forms in this extremal case.
Thus, at the small Planck scale, quantum effects would emerge. This implies that, instead of the temperature diverging as in the vacuum case, the matter sources proposed in this work lead the temperature to decrease until a black hole remnant is reached at $T=0$ and $r_{+}=r_{\text{ext}}$.
@@
Figure 11: This figure displays the values of the temperature corresponding to different values of the event horizon radius for $d=6$, $n=2$, $a=1$ and $l=10$
VI.2 Heat Capacity and radial evolution
We use the standard definition of heat capacity:
$C=\left(\frac{dT}{dM}\right)^{-1}=\frac{dM}{dr_{+}}\left(\frac{dT}{dr_{+}}\right)^{-1}$ (53)
In this section, we are interested in identifying the regions where the heat capacity is positive (stable), negative (unstable), and where it diverges, indicating the presence of phase transitions. In this connection, from Figure 1, the derivative $\frac{dM}{dr_{+}}$ is always positive, i.e., $\frac{dM}{dr_{+}}\geq 0$.
@@ VI.2 Heat Capacity and radial evolution
Since the sign of the derivative $\frac{dT}{dM}=\left(\frac{dT}{dr_{+}}\right)\left(\frac{dM}{dr_{+}}\right)^{-1}$ depends only on the sign of the derivative $\frac{dT}{dr_{+}}$, it follows that the sign of the heat capacity also depends solely on this latter derivative. In connection with the previous paragraph, from Figure 11, we can observe that, for our solution, the temperature exhibits three regions from left to right: i) $\frac{dT}{dr_{+}}>0$ due to the influence of matter sources at small scales; ii) and iii) which coincide with the vacuum LnFDGS AdS case, where $\frac{dT}{dr_{+}}<0$ and $\frac{dT}{dr_{+}}>0$, respectively. Thus, in our solution, there would be two phase transitions (analyzed from right to left): i) $C>0$ where $\frac{dT}{dM}>0$; ii) $C<0$ where $\frac{dT}{dM}<0$ (these two coincide with the vacuum LnFDGS AdS case, where both solutions become indistinguishable); iii) $C>0$ where $\frac{dT}{dM}>0$.
@@
Once both the vacuum solution and our regular black hole (RBH) become distinguishable, the heat capacities of the two solutions also differ at small scales. In this way, we can observe that: Moving from right to left, while the heat capacity of the vacuum AdS solution remains negative, implying that complete evaporation, as the parameter $M$ decreases toward $M\to 0$, would cause the temperature to diverge to infinity, in our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that, as the mass decreases to reach the extremal value $M=M_{\text{min}}$, the temperature decreases until a black hole remnant is reached at $T=0$, at which point the evaporation process halts. As mentioned earlier, this could occur at scales close to the Planck scale. That is, the nature of our matter sources, along with the avoidance of singularity formation, would lead to the existence of a remnant.
@@ VII DISCUSSION AND SUMMARIZE
As mentioned in the introduction, a model was recently proposed in Ref. Estrada and Aros (2025) to suppress physical singularities in Pure Lovelock (PL) gravity, where the energy density encodes the gravitational information of the vacuum solution through the Kretschmann scalar. However, in the PL framework, the inclusion of a negative cosmological constant is unfeasible, as its presence leads to the appearance of a curvature singularity in the spherically symmetric case.
To address the aforementioned issue, namely, the introduction of an energy density analogous to that in Ref. Estrada and Aros (2025) within Lovelock theories under the presence of a negative cosmological constant, in this work, instead of considering the PL action, which includes only a single term in the Lovelock series, we adopt the definition of the coupling constants given in Refs. Crisostomo et al. (2000); Aros et al. (2001). By including all terms in the Lovelock series up to order $n=\left[d/2\right]$, the resulting equations of motion possess an $n$-fold degenerate AdS ground state. In this way, the inclusion of a negative cosmological constant does not lead to the emergence of curvature singularities.
@@
The obtained solution is such that, for a mass parameter value slightly greater than $M_{\text{min}}$ (associated with an extremal radius), our RBH solution LnFDGS becomes indistinguishable from its vacuum counterpart LnFDGS Crisostomo et al. (2000); Aros et al. (2001) inside the event horizon, from a radial value $r=r_{-}<r_{+}$ up to infinity. If $r_{*}$ is of the order of the Planck scale, the geometric differences between both solutions, particularly the suppression of the singularity, would occur at quantum scales.
On the other hand, due to the high theoretical interest that the study of black holes in Lovelock gravity has garnered, it is natural to explore the properties of their shadows. However, in most solutions within this theory—such as our case of interest, the LnFDGS AdS solution—the analytical study of these shadows becomes challenging. In this regard, related to the study of shadows, a method has been proposed to obtain numerical graphical relationships between the event horizon, the photon sphere radius, and the shadow size in our case study. The case $n=2$, $d=7$ has been used as an example to explain our procedure. However, it is straightforward to verify that our proposal is applicable to other values of $n$ and $d$. We have also presented the graphical behavior for the cases $n=2$, $d=6,7,8$ and $n=3$, $d=8,9,10$. We note that, when using a common event horizon value as a reference for the same value of $n$, the photon sphere radius decreases as the number of dimensions increases. Additionally, when using a common photon sphere radius value as a reference for the same value of $n$, the size of the black hole shadow radius tends to decrease as the number of dimensions increases. It is also important to mention that, in all the cases studied numerically, the condition $r_{+}<r_{\text{sp}}<r_{\text{sh}}$ is satisfied.
@@
Regarding radial evolution and heat capacity, once the heat capacity of both the vacuum solution and our regular solution is distinguible at short scales, we observe a key difference: while the heat capacity of the vacuum AdS solution remains negative, implying that complete evaporation, as the parameter $M$ decreases toward $M\to 0$, would cause the temperature to diverge to infinity. In our regular AdS solution, the matter sources induce a new phase transition such that the heat capacity changes from negative to positive. This implies that as the mass decreases and approaches the extremal value $M=M_{\text{min}}$, the temperature also decreases until a black hole remnant is reached at $T=0$, at which point the evaporation process halts. As mentioned earlier, this could occur at scales close to the Planck scale. That is, the nature of our matter sources, along with the avoidance of singularity formation, would lead to the existence of a remnant.
@@ Acknowledgements
This work of RA was partially funded through FONDECYT-Chile 1220335. Milko Estrada is funded by the FONDECYT Iniciación Grant 11230247.